OpenStudy (he66666):

How do you solve the integral int_{0}^{pi/2} dx/(2 + cosx) ?

OpenStudy (he66666):

Sorry I'll type again

OpenStudy (he66666):

\[\int\limits_{0}^{\pi/2} dx/(2 + cosx) \]

OpenStudy (he66666):

the answer is \[\pi/(3\sqrt{3})\]

OpenStudy (he66666):

I tried partial integration by u = cosx + 2 and du = -sinxdx but it didn't work out..

OpenStudy (anonymous):

there's a tricky substitution that pretty much nobody who hasn't seen the integral before would see, but your substitution will do just fine. after you do that substitution (i'm ignoring the bounds for now): int -(sin(x)*u)^-1, but sin(x) = sqrt(1 -cos^2(x)) = sqrt(1 - (u -2)^2). which completely transforms the integral, try giving that a go

OpenStudy (he66666):

I converted the expression to the way you explained but I only got \[\int\limits_{}^{}1/(u \sqrt{1-(u-2)^2)}du\] and I don't know what to do now

OpenStudy (anonymous):

just to tidy things up you could do another substitution z = u - 2 -> dz = du, u = z + 2 so int((z+2)sqrt(1 - z^2))^-1 dz, which you could try to do by parts, i'll check that though

OpenStudy (he66666):

Weirdly, I got pi as the answer...

OpenStudy (anonymous):

Gaah, openstudy is pissing me off. I will try to write in latex, hang on.

OpenStudy (anonymous):

Again, I am unsure if this will be comprehensible. Sorry for not trying to clean this up :-) But, yeah, I think it's solvable now.

OpenStudy (he66666):

Thank you!